what is the quotient of a number and -9 plus four is -5.
Answer by ewatrrr(24785)


Hi
Let x represent the number
Question states***Write as we Read
n/-9 + 4 = -5
n - 36 = 45
n = 81
CHECKING our Answer***
81/-9 + 4 = -9 + 4 = -5